# A Mixture-of-Experts Framework with Log-Logistic Components for Survival Analysis on Histopathology Images

### v1

@@ 3.2 Graph-Guided Clustering (GGC)
Based on $S$, a $k$-nearest-neighbor (k-NN) graph is constructed, and patches are clustered into $K$ groups $L_{1},\dots,L_{K}$} using GPU-accelerated K-Means, guided by the graph structure. Each cluster represents a local tissue microenvironment such as tumor epithelium, stroma, or necrosis, promoting a biologically meaningful and spatially coherent organization across the slide.
@@ 3.3 Hierarchical Context Attention (HCA)
Intra-cluster representations are concatenated as $\widetilde{P}=\mathrm{Concat}(L^{\prime}_{k})_{k=1}^{K}$. A broadcast descriptor $R^{\prime}_{\mathrm{exp}}$, derived by averaging $R^{\prime}$ across clusters and expanding to match $\widetilde{P}$, is added residually to form $\widehat{P}$. Finally, $\widehat{P}$ is concatenated with remaining irrelevant patch features to yield $\mathcal{P}_{\text{final}}$.
@@ 3.4 Expert-Driven Mixture of Log-logistics
An Expert-Driven Mixture of Log-logistics module (Figure 1b) is employed to estimate continuous survival probabilities by modeling the full survival time distribution. Built on the mixture-of-experts framework [bishop2006pattern], it uses multiple experts to form Log-logistics Distributions(LLD) that capture heterogeneous survival patterns. The log-logistic probability density is given by
@@
The survival time $t$ of a whole slide image is directly modeled on its features derived via the attention pooling layer, as shown in Fig. 1(b). Each expert $e$ in the mixture models the conditional distribution over $t$ using a log-logistic mixture model (LLD):
@@
where each $k$ component follows a log-logistic distribution with scale $\alpha^{(e)}_{k}$ and shape $\beta^{(e)}_{k}$, while mixture weights $\lambda^{(e)}_{k}$ are computed using a softmax over the $K$ components. Learnable global anchors $(P_{\alpha},P_{\beta})\in\mathrm{R}^{K}$, corresponding to scale and shape parameters, are linearly projected for each expert:
$\alpha^{(e)}=W^{(e)}_{\alpha}P_{\alpha},\qquad\beta^{(e)}=\text{softplus}\!\big(W^{(e)}_{\beta}P_{\beta}\big),$ (12)
@@ 3.5 Training Objective and Regularization
Table 1: Left: TCGA performance (TDC, higher is better). Right: Ablations on $q$ in QGPS, GGC, HCA. Best TDC results in bold.
@@ 4.2 Training Configuration
The quantile hyperparameter is set to $q=0.25$, retaining the top $75\%$ of patches based on logits. Each patch in the k-NN graph connects to its 10 nearest neighbors, which are clustered into certain groups and processed by hierarchical attention with 8 heads. The Expert-Driven Mixture of Log-Logistics module employs five experts, each modeling a log-logistic distribution (LLD) with $K=100$ components. Model is trained for 20 epochs with Adam (lr = $2\times 10^{-4}$, weight decay = $1\times 10^{-3}$, dropout = 0.1) and a batch size of 1. We report mean $\pm$ standard deviation of the Time-Dependent Concordance Index (TDC) [han2022survivalmixturedensitynetworks] metric over 5-fold cross-validation. All experiments are conducted on an NVIDIA A6000 GPU.
@@ 5.2 Ablation Studies and Interpretability
To assess each module’s contribution, ablation studies were performed on the Quantile-Gated Patch Selection, Graph-Guided Clustering, and Hierarchical Context Attention components (Table 1(b)). Varying the quantile parameter showed that q=0.25q{=}0.25 offers the best trade-off, retaining prognostic regions while filtering irrelevant tissue. Removing any module causes a noticeable drop in TDC, confirming their complementary roles in denoising and spatial reasoning. For interpretability, patients were divided into high- and low-risk groups using predicted survival scores, and Kaplan–Meier (KM) curves were plotted (Fig. 2). The groups show distinct survival trends with statistically significant separation (log-rank p=0.046p{=}0.046 for LUAD, p=0.005p{=}0.005 for KIRC, p=0.017p{=}0.017 for BRCA), validating that the expert-guided architecture captures prognostically relevant morphology. Furthermore, dynamic patch selection and graph-guided attention improve localization of high-risk regions, enhancing both predictive accuracy and clinical interpretability.
@@ 6 Conclusion
We proposed a unified framework for WSI-based survival prediction integrating quantile-based patch selection, graph-guided clustering, hierarchical context attention, and expert-driven mixture of log-logistics. By modeling local–global tissue interactions through clustered attention and estimating survival distributions via multiple experts, the method improves calibration and discrimination. Experiments on TCGA-LUAD, TCGA-KIRC, and TCGA-BRCA show consistent gains in time-dependent concordance over pathology-based and multimodal baselines. Future work will extend the framework toward multimodal and uncertainty-aware patient-wise survival prediction for clinical reliability.

### v2

@@ 3.2 Graph-Guided Clustering (GGC)
Based on $S$, a $k$-nearest-neighbor (k-NN) graph is constructed, and patches are then clustered into $G$ groups $L_{1},\dots,L_{G}$} using GPU-accelerated K-Means, guided by the graph structure. Each cluster represents a local tissue microenvironment such as tumor epithelium, stroma, or necrosis, promoting a biologically meaningful and spatially coherent organization across the slide.
@@ 3.3 Hierarchical Context Attention (HCA)
Intra-cluster representations are concatenated as $\widetilde{P}=\mathrm{Concat}(L^{\prime}_{g})_{g=1}^{G}$. A broadcast descriptor $R^{\prime}_{\mathrm{exp}}$, derived by averaging $R^{\prime}$ across clusters and expanding to match $\widetilde{P}$, is added residually to form $\widehat{P}$. Finally, $\widehat{P}$ is concatenated with remaining irrelevant patch features to yield $\mathcal{P}_{\text{final}}$.
@@ 3.4 Expert-Driven Mixture of Log-logistics
An Expert-Driven Mixture of Log-Logistics module (Figure 1b) models survival probabilities, with each expert specializing in distinct WSI features $\mathbf{z}_{\text{WSI}}$. Built on the mixture-of-experts framework [bishop2006pattern], it uses multiple experts to form Log-logistics Distributions(LLD) that capture heterogeneous survival patterns. The log-logistic probability density is given by
@@
The survival time $t$ of a whole slide image is directly modeled on its features $\mathbf{z}_{\text{WSI}}$ derived via the attention pooling layer, as shown in Fig. 1(b). Each expert $e$ in the mixture models the conditional distribution over $t$ using a log-logistic mixture model (LLD):
@@
where each $k$ component follows a log-logistic distribution with scale $\alpha^{(e)}_{k}$ and shape $\beta^{(e)}_{k}$, while mixture weights $\lambda^{(e)}_{k}$ are computed using a softmax over the $K$ components. Learnable global anchors $(P_{\alpha},P_{\beta})\in\mathrm{R}^{K}$, corresponding to scale and shape parameters, are projected for each expert as below:
$\alpha^{(e)}=\text{softplus}\!\big(W^{(e)}_{\alpha}P_{\alpha}\big),\quad\beta^{(e)}=\text{softplus}\!\big(W^{(e)}_{\beta}P_{\beta}\big)$ (12)
@@ 3.5 Training Objective and Regularization
Table 1: Left: TCGA performance (TDC, higher is better). Right: Ablations on $q$ in QGPS, GGC, HCA and $e$ in EDMLL. Best TDC results in bold.
@@ 4.2 Training Configuration
The quantile hyperparameter is set to $q=0.25$, retaining the top $75\%$ of patches based on logits. The $k$-NN graph constructed from the similarity matrix $S$ connects each point to its 10 nearest neighbors. Based on the graph connectivity, the patch indices are clustered into $G$ groups of 64 each, and these grouped patches are subsequently processed by an 8-head MHSA within the HCA module. The Expert-Driven Mixture of Log-Logistics module employs five experts, each modeling a log-logistic distribution (LLD) with $K=100$ components. Model is trained for 20 epochs with Adam (lr = $2\times 10^{-4}$, weight decay = $1\times 10^{-3}$, dropout = 0.1) and a batch size of 1. Results are reported as mean $\pm$ standard deviation of the Time-Dependent Concordance Index (TDC) [han2022survivalmixturedensitynetworks] metric over 5-fold cross-validation. All experiments are conducted on an NVIDIA A6000 GPU.
@@ 5.2 Ablation Studies and Interpretability
Ablation studies are conducted to evaluate the contributions of Quantile-Gated Patch Selection, Graph-Guided Clustering, Hierarchical Context Attention and Expert-Driven Mixture of Log-logistics(Table 1(b)). Varying the quantile parameter showed that q=0.25q{=}0.25 provided the best trade-off, preserving prognostic regions while filtering irrelevant tissue. Eliminating any module resulted in a reduction in TDC, confirming their contributions to noise suppression and contextual feature integration. Single-expert model fell short, whereas five experts yielded superior results, underscoring the benefit of multi-expert modeling. For interpretability, patients were categorized into high and low-risk groups based on their predicted survival scores. The Kaplan–Meier curves(Fig. 2) revealed distinct survival trends with significant separation (log-rank p=0.046p{=}0.046 for LUAD, p=0.005p{=}0.005 for KIRC, p=0.017p{=}0.017 for BRCA), confirming that the expert-driven architecture effectively captures prognostically relevant morphological patterns.
@@ 6 Conclusion
We proposed a unified framework for WSI-based survival prediction integrating quantile-based patch selection, graph-guided clustering, hierarchical context attention, and expert-driven mixture of log-logistics. By modeling local–global tissue interactions through clustered attention and estimating survival distributions via multiple experts, the method improves calibration and discrimination. Experiments on TCGA-LUAD, TCGA-KIRC, and TCGA-BRCA show consistent gains in time-dependent concordance over pathology-based and multimodal baselines. Future work will extend the framework towards multimodal and uncertainty-aware patient-wise survival prediction for clinical reliability.